The larger of two numbers is 15 more then twice the smaller number. if the sum of the numbers is 60 find the numbers. put the smaller number first and the larger number second.
Answer by ikleyn(52778)

.

Let x be the larger number, y be the smaller number.


From the condition, we have these two equations


    x = 2y + 15     (1)

    x + y = 60      (2)


From equation (1), substitute this expression x = 2y + 15  into the second equation, replacing x there.


You will get a single equation then, for only one unknown y


    (2y + 15) + y = 60.


Simplify and find y


    3y = 60 - 15

    3y = 45

     y = 45/3 = 15.


Then from equation (1),  x = 2y+15 = 2*15 + 15 = 45.


ANSWER.  The larger number is 45;  the smaller number is 15.

Solved.

You can check it on your own that my answer SATISFIES the given equations.